Vic can bind 200 booklets in 1 hr..Annie can do the same job in 50 minutes.How long will it take them to finish if Vic binds the first 50 booklets, then annie binds 30minutes,then Vic finishes th job? *Thank YOu *
Answer by mananth(16946)

Vic can bind 200 booklets in 1 hr.
50 booklets he will bind in 1/4 hr.
..
annie binds 200 booklets in 50 minutes
so in 1/2 hour she binds 120 booklets.
..
170 booklets have been bound in 1/4 +1/2 hour=3/4 hour
balance 30 booklets vic binds.
vic bind 200 booklets in 1 hour
so 30 booklets he will do in 30/200=3/20 hour
..
total time = 3/4 +3/20=15+3 /20=18/20 hours
= 9/10 hours.